# Guiding-center dynamics in a screw-pinch magnetic field

### v1

@@ I.1 Adiabatic invariance of the magnetic moment
Here, the dimensionless parameter $\epsilon$ appears as a result of the renormalization of the particle’s mass $m\rightarrow\epsilon\,m$ as a way to introduce the dimensional mass-over-charge ratioNorthrop (1963) $m/e$ as an expansion parameter, while the guiding-center canonical momentum ${\bf P}_{\rm gc}\equiv{\sf T}_{\rm gc}^{-1}{\bf P}$ is defined as the guiding-center push-forward of the canonical momentum ${\bf P}\equiv(e/c)\,{\bf A}({\bf x})+\epsilon\,m{\bf v}$, and $\epsilon\,\mbox{$\rho$}_{\rm gc}\equiv{\sf T}_{\rm gc}^{-1}{\bf x}-{\bf X}_{\rm gc}$ is the guiding-center gyroradius, which is defined as the difference between the guiding-center push-forward of the particle position $\bf x$} and the guding-center position ${\bf X}_{\rm gc}$. In addition, the Lie-transform guiding-center push-forward operatorLittlejohn (1982) ${\sf T}_{\rm gc}^{-1}$ involved in Eq. (1) is defined for an arbitrary phase-space function $f$ by the series expansion in powers of $\epsilon$:
@@
where the Lie-transform operator ${\sf T}_{n}^{-1}\equiv\exp(\epsilon^{n}{\mathcal{L}}_{n})$ is defined in terms of the $n$th-order Lie derivative ${\mathcal{L}}_{n}\equiv{\sf G}_{n}\cdot{\sf d}$ generated by the vector field ${\sf G}_{n}$, whose purpose is the remove the gyroangle dependence from the guiding-center dynamics. The expansion of the action integral (1) in powers of $\epsilon$ yields the so-called Kruskal asymptotic seriesKruskal (1962) for the guiding-center magnetic-moment adiabatic invariant.
While the integral definition (1) for the guiding-center magnetic moment $\mu_{\rm gc}$ yields the correct expression for the lowest-order magnetic moment $\mu_{0}\equiv\epsilon^{2}\,m|{\bf v}_{\bot}|^{2}/(2\,eB/m)$, it is rarely used to obtain higher-order terms because it involves multiple perturbation expansions in powers of $\epsilon$ before performing the gyroangle averages needed to obtain an explicit expression at a desired perturbation order. Instead, by using the standard guiding-center perturbation theoryLittlejohn (1983); Cary and Brizard (2009) associated with the phase-space transformation from particle coordinates $({\bf x},{\bf p}=m{\bf v})$ to guiding-center coordinates $({\bf X}_{\rm gc},P_{\|{\rm gc}},\mu_{\rm gc},\zeta_{\rm gc})$, the guiding-center magnetic moment is defined as an explicit asymptotic expansion in powers of $\epsilon$:
$\mu_{\rm gc}\;\equiv\;\mu_{0}\;+\;\epsilon\,\mu_{1}\;+\;\epsilon^{2}\,\mu_{2}\;+\;\cdots\;=\;\mu_{0}\;+\;\epsilon\,G_{1}^{\mu}\;+\;\epsilon^{2}\left(G_{2}^{\mu}\;+\;\frac{1}{2}\,{\sf G}_{1}\cdot{\sf d}G_{1}^{\mu}\right)\;+\;\cdots,$ (3)
where the first-order correction $\mu_{1}\equiv G_{1}^{\mu}$ is associated with first-order magnetic-field non-uniformity, while the terms $\mu_{n}$ ($n\geq 2)$ involve higher-order correctionsBurby et al. (2013); Tronko and Brizard (2015). Unfortunately, there are concerns about the validity of the asymptotic expansion (3) when the dimensionless parameter $\epsilon$ is not small, either because of steep gradientsBrizard (2017) or when considering energetic particlesJ. W. Burby et al. (2025).
@@ I.2 Kruskal Identity
The purpose of the present paper is to investigate the non-perturbative definition of the guiding-center magnetic moment in the case of doubly-symmetric time-independent magnetic geometries, where the two ignorable spatial coordinates $(q^{1},q^{2})$ are associated with their respective invariant canonical momenta $(P_{1},P_{2})$. In this particular case, the Lagrangian dynamics takes place in a reduced two-dimensional space $(q_{3},dq_{3}/dt)$, where the reduced Lagrangian is derived by the Routh reduction procedureBrizard (2015)
@@
$P_{3}$ $\equiv$ $\partial L_{r}/\partial(dq^{3}/dt)\;=\;m\,dq_{3}/dt,$
@@
there exists a reduced action integral
@@
$J_{3}(E,P_{1},P_{2})\;=\;\frac{1}{2\pi}\;\int_{0}^{2\pi}P_{3}\,\frac{\partial q^{3}}{\partial\zeta}\;d\zeta\;\equiv\;\mu_{\rm gc},$ (6)
where the ordering parameter $\epsilon$ appears in a non-perturbative setting in the definition of $J_{3}$, which yields a non-perturbative definition of the magnetic moment $\mu_{\rm gc}$.
In recent papers J. W. Burby et al. (2025); Holas et al. (2025), the Kruskal identity (6) was proved for two doubly-symmetric magnetic geometries. First, this identity was proved up to fourth order in $\epsilon$ (including second-order effects in magnetic field non-uniformity) for the case of a straight magnetic field $^{\bf B}=(1+y)\,\widehat{\sf z}$ with a constant gradient, where the canonical momenta $(P_{x},P_{z})$ in the ignorable directions $(x,z)$ are invariant. Second, a screw-pinch magnetic field was considered, with the magnetic field expressed in terms of an Eulerian representation
@@
in cylindrical geometry, where the axial magnetic flux $\psi(r)$ and the azimuthal magnetic field $\Psi(\psi)$ are assumed to be functions of the radial position $r$ only, and the canonical momenta $(P_{\theta},P_{z})$ are constants of motion. Unfortunately, the proof of Kruskal’s identity (6) for the screw-pinch magnetic field (7) required considerable use of computer algebra up to third order in $\epsilon$ (i.e., including first-order corrections in magnetic non-uniformity).
@@ I.3 Organization of the paper
$\widehat{{\sf b}}\;=\;\cos\Theta\;\widehat{\sf z}\;+\;\sin\Theta\;\widehat{\theta}$, (9)
@@
For the purpose of explicitly proving the Kruskal identity (6) for a doubly-symmetric screw-pinch magnetic geometry, we begin with the Lie-transform perturbation expansion (3) of the guiding-center magnetic moment in Sec. V. Here, the guiding-center transformation for the particle motion in screw-pinch magnetic geometry, which is reviewed in App. A, is used to obtain an explicit expression for the guiding-center magnetic moment $\mu_{\rm gc}$ up to first order in magnetic non-uniformity (i.e., third order in the ordering parameter $\epsilon$). In Sec. VI, we complete the proof of the Kruskal identity (6) by expanding the reduced radial action $J_{r}(E;P_{\theta},P_{z})$ in powers of $\epsilon$ and perform explicit gyroangle averages to show that $J_{r}\equiv\mu_{\rm gc}$ up to third order in $\epsilon$. Hence, amazingly, the guiding-center magnetic moment $\mu_{\rm gc}=J_{r}$ is defined as an exact radial action invariant that depends non-perturbatively on the ordering parameter $\epsilon$. Finally, in Sec. VII, we discuss how the guiding-center transformation used in Secs. V-VI can be extended to fourth order in $\epsilon$ (i.e., second order in magnetic-field non-uniformity).
@@ II Screw Pinch Magnetic Geometry
We consider a screw-pinch magnetic field (8) in cylindrical geometry, where we will use the radial pitch angle $\Theta(r)=\pi\,(r/a)^{2}/2\leq\pi/2$ for explicit results in what follows, and we will assume that the plasma is confined in the region $r\leq a$ (i.e., $r=a$ denotes the edge of the plasma).
@@
$\frac{d\widehat{{\sf b}}}{ds}$ $=$ $\widehat{{\sf b}}\,\mbox{$\cdot$}\,\nabla\widehat{{\sf b}}\;=\;\kappa\;\widehat{\sf e}_{1},$ (10)
$\frac{d\widehat{\sf e}_{1}}{ds}$ $=$ $\widehat{{\sf b}}\,\mbox{$\cdot$}\,\nabla\widehat{\sf e}_{1}\;=\;-\,\kappa\,\widehat{{\sf b}}\;+\;\tau\;\widehat{\sf e}_{2},$ (11)
$\frac{d\widehat{\sf e}_{2}}{ds}$ $=$ $\widehat{{\sf b}}\,\mbox{$\cdot$}\,\nabla\widehat{\sf e}_{2}\;=\;-\,\tau\;\widehat{\sf e}_{1},$ (12)
@@
$\left(\kappa,\frac{}{}\tau\right)\;\equiv\;\left(\frac{1}{r}\,\sin^{2}\,\Theta,\;\frac{1}{r}\,\sin\,\Theta\,\cos\,\Theta\right).$ (14)
@@
in terms of the Frenet-Serret torsion $\tau$ and the magnetic-pitch shear $\Theta^{\prime}(r)$.
@@
$\equiv$ $\tau\,\widehat{{\sf b}}\;-\;\left(\frac{1}{r}-\kappa\right)\;\widehat{\sf e}_{2}\;\equiv\;\tau\,\widehat{{\sf b}}\;-\;R_{\bot}\,\widehat{\sf e}_{2},$
@@ III Geometric Newtonian Mechanics
where $\overline{\bf B}\equiv{\bf B}/B_{0}$ is the normalized magnetic field, and $\epsilon\equiv(mv_{0}/eB_{0})/L_{0}$ is the ratio of the characteristic gyroradius $(mv_{0}/eB_{0})$ to the magnetic-nonuniformity length scale $L_{0}$. In what follows, we will use the notation $(t,{\bf r},{\bf v})$, without the overbar, to denote the normalized time, particle position, and particle velocity.
@@ III.1 Velocity representation in screw-pinch magnetic geometry
$\equiv$ $\left(r\,\frac{d\theta}{dt}\,\sin\,\Theta\;+\;\frac{dz}{dt}\,\cos\,\Theta\right)\widehat{{\sf b}}\;-\;\frac{dr}{dt}\,\widehat{\sf e}_{1}\;+\;\left(\frac{dz}{dt}\,\sin\,\Theta\;-\;r\,\frac{d\theta}{dt}\,\cos\,\Theta\right)\widehat{\sf e}_{2},$
where the normalized screw-pinch pitch angle is $\Theta=\pi\,r^{2}/2$. We may also express the particle velocity (III.1) as $^{\bf v}\equiv v_{\|}\,\widehat{{\sf b}}+v_{\bot}\,\widehat{\bot}$ in terms of the rotating unit vectors $\widehat{\bot}\equiv\widehat{\rho}\,\mbox{$\times$}\,\widehat{{\sf b}}$, where
$\widehat{\bot} \displaystyle\equiv$ $\widehat{\rho}\,\mbox{$\times$}\,\widehat{{\sf b}}\;=\;\frac{\partial\widehat{\rho}}{\partial\zeta}\;=\;-\,\sin\zeta\;\widehat{\sf e}_{1}\;-\;\cos\zeta\;\widehat{\sf e}_{2},$ (20)
$\widehat{\rho} \displaystyle\equiv$ $\widehat{{\sf b}}\,\mbox{$\times$}\,\widehat{\bot}\;=\;-\,\frac{\partial\widehat{\bot}}{\partial\zeta}\;=\;\cos\zeta\;\widehat{\sf e}_{1}\;-\;\sin\zeta\;\widehat{\sf e}_{2},$ (21)
@@
Here, the cylindrical velocity components (24) and (26)-(27) are expressed in screw-pinch magnetic geometry in terms of the pitch angle $\Theta$ and the velocity components $(v_{\|},v_{\bot},\zeta)$.

### v2

@@ I.1 Adiabatic invariance of the magnetic moment
Here, the dimensionless parameter $\epsilon$ appears as a result of the renormalization of the particle’s mass $m\rightarrow\epsilon\,m$ as a way to introduce the dimensional mass-over-charge ratioNorthrop (1963) $m/e$ as an expansion parameter, while the guiding-center canonical momentum ${\bf P}_{\rm gc}\equiv{\sf T}_{\rm gc}^{-1}{\bf P}$ is defined as the guiding-center push-forward of the canonical momentum ${\bf P}\equiv(e/c)\,{\bf A}({\bf x})+\epsilon\,m{\bf v}$, and $\epsilon\,\mbox{$\rho$}_{\rm gc}\equiv{\sf T}_{\rm gc}^{-1}{\bf x}-{\bf X}_{\rm gc}$ is the guiding-center gyroradius, which is defined as the difference between the guiding-center push-forward of the particle position $\bf x$} and the guding-center position ${\bf X}_{\rm gc}$. In addition, the Lie-transform guiding-center push-forward operatorLittlejohn (1982a) ${\sf T}_{\rm gc}^{-1}$ involved in Eq. (1) is defined for an arbitrary phase-space function $f$ by the series expansion in powers of $\epsilon$:
@@
where the Lie-transform operator ${\sf T}_{n}^{-1}\equiv\exp(-\epsilon^{n}{\mathcal{L}}_{n})$ is defined in terms of the $n$th-order Lie derivative ${\mathcal{L}}_{n}\equiv{\sf G}_{n}\cdot{\sf d}$ generated by the vector field ${\sf G}_{n}$, whose purpose is to remove the gyroangle dependence from the guiding-center dynamics. The expansion of the action integral (1) in powers of $\epsilon$ yields the so-called Kruskal asymptotic seriesKruskal (1962) for the guiding-center magnetic-moment adiabatic invariant.
While the integral definition (1) for the guiding-center magnetic moment $\mu_{\rm gc}$ yields the correct expression for the lowest-order magnetic moment $\mu_{0}({\bf x},{\bf v})\equiv\epsilon^{2}\,m|{\bf v}_{\bot}|^{2}/(2\,eB({\bf x})/m)$, which is written here with units of action, it is rarely used to obtain higher-order terms because it involves multiple perturbation expansions in powers of $\epsilon$ before performing the gyroangle averages needed to obtain an explicit expression at a desired perturbation order. Instead, by using the standard guiding-center perturbation theoryLittlejohn (1983); Cary and Brizard (2009) associated with the phase-space transformation from particle coordinates $({\bf x},{\bf p}=m{\bf v})$ to guiding-center coordinates $({\bf X}_{\rm gc},P_{\|{\rm gc}},\mu_{\rm gc},\zeta_{\rm gc})$, the guiding-center magnetic moment is defined as an explicit asymptotic expansion in powers of $\epsilon$:
$\mu_{\rm gc}({\bf x},{\bf v})\;\equiv\;\mu_{0}\;+\;\epsilon\,\mu_{1}\;+\;\epsilon^{2}\,\mu_{2}\;+\;\cdots\;=\;\mu_{0}\;+\;\epsilon\,G_{1}^{\mu}\;+\;\epsilon^{2}\left(G_{2}^{\mu}\;+\;\frac{1}{2}\,{\sf G}_{1}\cdot{\sf d}G_{1}^{\mu}\right)\;+\;\cdots,$ (3)
where the first-order correction $\mu_{1}\equiv G_{1}^{\mu}$ is associated with first-order magnetic-field non-uniformity, while the terms $\mu_{n}$ ($n\geq 2)$ involve higher-order correctionsBurby et al. (2013); Tronko and Brizard (2015). Unfortunately, there are concerns about the validity of the asymptotic expansion (3) when the dimensionless parameter $\epsilon$ is not small, either because of steep gradientsBrizard (2017) or when considering energetic particlesJ. W. Burby et al. (2025). It should be pointed out, however, that the comparison of full-orbit particle trajectories with guiding-center orbits depends on the faithfulness of the guiding-center equations, as was noted by Belova et al. Belova et al. (2003) and Brizard and Hodgeman Brizard and Hodgeman (2023). It has been shown by Brizard and Hodgeman Brizard and Hodgeman (2023), for example, that the traditional guiding-center formalism Littlejohn (1983) can be improved by retaining corrections associated with guiding-center polarization Tronko and Brizard (2015), which were not included in the guiding-center studies cited by Burby et al. J. W. Burby et al. (2025).
A more important comment concerns the conundrum associated with the comparison the integral expression (1) with the local expression (3). On the one hand, the integral expression (1) is explicitly independent of the gyroangle while, on the other hand, the local expression (3) is explicitly dependent on the gyroangle. In the original work on non-perturbation magnetic-moment invarianceJ. W. Burby et al. (2025), where the magnetic field $^{\bf B}=B(x,y)\,\widehat{\sf z}$ has a single spatial symmetry axis, a Poincaré approach is adopted where the local expression (3) is restricted to the Poincaré section defined at the gyroangle $\zeta=0$ [see Eq. (4)J. W. Burby et al. (2025)], while the integral expression (1) is estimated by sophisticated statistical analysis of the full-orbit dynamics as orbits intersect the Poincaré section for two different values of $\epsilon$.
@@ I.2 Kruskal Identity
The purpose of the present paper is to investigate the non-perturbative definition of the guiding-center magnetic moment in the case of doubly-symmetric time-independent magnetic geometries, where the two ignorable spatial coordinates $(q^{1},q^{2})$ are associated with their respective invariant canonical momenta $(P_{1},P_{2})$. In this particular case, the Lagrangian dynamics takes place in a reduced two-dimensional space $(q^{3},dq^{3}/dt)$, where the reduced Lagrangian is derived by the Routh reduction procedureBrizard (2015)
@@
$P_{3}$ $\equiv$ $\partial L_{r}/\partial(dq^{3}/dt)\;=\;m\,dq^{3}/dt,$
@@
there exists a reduced action integralGoldstein et al. (2002)
@@
$J_{r}(E,P_{1},P_{2})\;=\;\frac{1}{2\pi}\;\int_{0}^{2\pi}P_{3}\,\frac{\partial q^{3}}{\partial\zeta}\;d\zeta\;\equiv\;\mu_{\rm gc},$ (6)
where the ordering parameter $\epsilon$ appears in a non-perturbative setting in the definition of $J_{r}$, which yields a non-perturbative definition of the magnetic moment $\mu_{\rm gc}$. We note that since the left-hand side of Eq. (6) is explicitly gyroangle-independent, the guiding-center magnetic moment (3), which is defined in particle phase space, must be gyroangle-independent as well when the full-orbit particle solution is inserted. More importantly, this gyroangle independence must occur without gyroangle-averaging. Hence, a second important constraint associated with the Kruskal identity (6) is that the expansion of the guiding-center magnetic moment (3) must be explicitly gyroangle-independent, which yields a consistency constraint between the full-orbit particle dynamics and the guiding-center transformation.
In a recent paper Holas et al. (2025), the Kruskal identity (6) was proved for two doubly-symmetric magnetic geometries. First, this identity was proved up to fourth order in $\epsilon$ (including second-order effects in magnetic field non-uniformity) for the case of a straight magnetic field $^{\bf B}=(1+y)\,\widehat{\sf z}$ with a constant gradientBrizard (2017, 2022), where the canonical momenta $(P_{x},P_{z})$ in the ignorable directions $(x,z)$ are invariant. This case is briefly reviewed in App. A, where the Kruskal identity (6) is proved and the constraint that the gyroangle independence of the guiding-center magnetic moment (3) is verified.
Second, a screw-pinch magnetic field was considered Holas et al. (2025), with the magnetic field expressed in terms of an Eulerian representation
@@
in cylindrical geometry, where the axial magnetic flux $\psi(r)$ and the azimuthal magnetic flux $\Psi(\psi)$ are assumed to be functions of the radial position $r$ only, and the canonical momenta $(P_{\theta},P_{z})$ are constants of motion. Unfortunately, the proof of Kruskal’s identity (6) for the screw-pinch magnetic field (7) required considerable use of computer algebra up to third order in $\epsilon$ (i.e., including first-order corrections in magnetic non-uniformity). The main problem associated with an explicit proof of the Kruskal identity (6) for this case is that the solution for the azimuthal magnetic flux $\Psi(\psi)$ cannot be obtained in closed form, which prevented the Lagrangian formalism from producing explicit results. The main purpose of the present work is to show that, in contrast to the Lagrangian formalism adopted by Holas et al.Holas et al. (2025), the adoption of the equivalent Newtonian formalism will yield an explicit proof of the Kruskal identity (6), and show that the expansion of the guiding-center magnetic moment (3) is explicitly gyroangle-independent.
@@ I.3 Organization of the paper
$\widehat{{\sf b}}(r,\theta)\;=\;\cos\Theta(r)\;\widehat{\sf z}\;+\;\sin\Theta(r)\;\widehat{\theta}(\theta),$ (9)
@@
For the purpose of explicitly proving the Kruskal identity (6) for a doubly-symmetric screw-pinch magnetic geometry, we begin with the Lie-transform perturbation expansion (3) of the guiding-center magnetic moment in Sec. V. Here, the guiding-center transformation for the particle motion in screw-pinch magnetic geometry, which is reviewed in App. B, is used to obtain an explicit expression for the guiding-center magnetic moment $\mu_{\rm gc}$ up to first order in magnetic non-uniformity (i.e., third order in the ordering parameter $\epsilon$). In Sec. VI, we complete the proof of the Kruskal identity (6) by expanding the reduced radial action $J_{r}(E;P_{\theta},P_{z})$ in powers of $\epsilon$ and perform explicit gyroangle averages to show that $J_{r}\equiv\mu_{\rm gc}$ up to third order in $\epsilon$. Hence, as expected, the guiding-center magnetic moment $\mu_{\rm gc}=J_{r}$ is defined as an exact radial action invariant that depends non-perturbatively on the ordering parameter $\epsilon$. Finally, in Sec. VII, we discuss how the guiding-center transformation used in Secs. V-VI can be extended to fourth order in $\epsilon$ (i.e., second order in magnetic-field non-uniformity).
@@ II Screw Pinch Magnetic Geometry
$\frac{d\widehat{{\sf b}}}{ds}$ $=$ $\frac{\sin\Theta}{r}\;\frac{\partial\widehat{{\sf b}}(r,\theta)}{\partial\theta}\;=\;\kappa(r)\;\widehat{\sf e}_{1},$ (10)
$\frac{d\widehat{\sf e}_{1}}{ds}$ $=$ $\frac{\sin\Theta}{r}\;\frac{\partial\widehat{\sf e}_{1}(r,\theta)}{\partial\theta}\;=\;-\,\kappa(r)\;\widehat{{\sf b}}\;+\;\tau(r)\;\widehat{\sf e}_{2},$ (11)
$\frac{d\widehat{\sf e}_{2}}{ds}$ $=$ $\frac{\sin\Theta}{r}\;\frac{\partial\widehat{\sf e}_{2}(r,\theta)}{\partial\theta}\;=\;-\,\tau(r)\;\widehat{\sf e}_{1},$ (12)
@@
$\left(\kappa,\;\tau\right)\;\equiv\;\left(\frac{1}{r}\,\sin^{2}\,\Theta,\;\frac{1}{r}\,\sin\,\Theta\,\cos\,\Theta\right).$ (14)
@@
in terms of the Frenet-Serret torsion $\tau(r)$ and the magnetic-pitch shear $\Theta^{\prime}(r)$.
@@
$\equiv$ $\tau\,\widehat{{\sf b}}\;-\;\left(\frac{1}{r}-\kappa\right)\;\widehat{\sf e}_{2}\;\equiv\;\tau(r)\,\widehat{{\sf b}}\;-\;R_{\bot}(r)\,\widehat{\sf e}_{2},$
@@ III Geometric Newtonian Mechanics
where $\overline{\bf B}\equiv{\bf B}/B_{0}$ is the normalized magnetic field, and $\epsilon\equiv(mv_{0}/eB_{0})/L_{0}$ is the ratio of the characteristic gyroradius $(mv_{0}/eB_{0})$ to the magnetic-nonuniformity length scale $L_{0}$. We note, here, that the characteristic range for $\epsilon$ is $\epsilon\simeq 2\,\times\,10^{-3}$ for a 10 keV proton and $\epsilon\simeq 0.08$ for a 3.5 MeV alpha particle in a 5T magnetic field with $L_{0}=100$ cm. In what follows, we shall simply assume that $\epsilon\ll 1$ can be used as a suitable perturbation parameter. In addition, we will use the notation $(t,{\bf r},{\bf v})$, without the overbar, to denote the normalized time, particle position, and particle velocity.
@@ III.1 Velocity representation in screw-pinch magnetic geometry
$\equiv$ $\left(r\,\frac{d\theta}{dt}\,\sin\,\Theta\;+\;\frac{dz}{dt}\,\cos\,\Theta\right)\widehat{{\sf b}}\;-\;\frac{dr}{dt}\,\widehat{\sf e}_{1}\;+\;\left(\frac{dz}{dt}\,\sin\,\Theta\;-\;r\,\frac{d\theta}{dt}\,\cos\,\Theta\right)\widehat{\sf e}_{2}.$
[…]
$\widehat{\bot} \displaystyle\equiv$ $\frac{\partial\widehat{\rho}}{\partial\zeta}\;=\;-\,\sin\zeta\;\widehat{\sf e}_{1}\;-\;\cos\zeta\;\widehat{\sf e}_{2},$ (20)
$\widehat{\rho} \displaystyle\equiv$ $-\,\frac{\partial\widehat{\bot}}{\partial\zeta}\;=\;\cos\zeta\;\widehat{\sf e}_{1}\;-\;\sin\zeta\;\widehat{\sf e}_{2},$ (21)
@@
Here, the cylindrical velocity components (24) and (26)-(27) are expressed in screw-pinch magnetic geometry in terms of the pitch angle $\Theta$ and the velocity coordinates $(v_{\|},v_{\bot},\zeta)$.